Find the least common multiple:
(a) 6 and 8


(b) 15 and 36

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

Find the least common multiple:
(a) 6 and 8
6 = 2*3
8 = 2^3
---
lcm has each prime factor in its highest power:
lcm = 3*2^3 = 24
------------------------------
(b) 15 and 36
15 = 3*5
36 = 2^2*3^2
----
lcm = 2^2*3^2*5 = 180
============================
Cheers,
Stan H.

Answer by edjones(8007)

Factor each number. To use this method factor each of the numbers into primes. Then for each different prime number in all of the factorizations, do the following...
Count the number of times each prime number appears in each of the factorizations.
For each prime number, take the largest of these counts.
Write down that prime number as many times as you counted for it in step 2.
The least common multiple is the product of all the prime numbers written down.
(From: help with fractions.com)
(b) 15 and 36
15: 3, 5
36: 2, 2, 3, 3
.
we need 2,2, 3, 3, 5
2*2*3*3*5=180
.
Ed